OpenStudy (anonymous):

The length of a rectangle is 3 centimeters less than twice its width. The perimeter of the rectangle is 48 cm. What are the dimensions of the rectangle? A)length = 15 cm; width = 9 cm (B)length = 11 cm; width = 7 cm (C)length = 12 cm; width = 12 cm (D)length = 14 cm; width = 10 cm

OpenStudy (anonymous):

answer is a) l = 2b -3 2b +2l = 48 solving we get l = 15 b = 9

OpenStudy (ajprincess):

Let the length of the rectangle be l and width be w Since it is given that the length is 3cm less than twice the width we have l=2w-3 But Perimeter P=2(w+l) 48=2(w+l) Dividing both sides by 2 we get 24=l+w 28=2w-3+w solving this we will get the value of w. Substituting in l=2w-3 we can find l

OpenStudy (ajprincess):

Does that help? @Drea1994

OpenStudy (anonymous):

yes it do thank yall soo much

OpenStudy (ajprincess):

Welcome:)